# Effective algorithms for tensor train decomposition via the UTV framework

### v1

@@ III-B TT-URV case
0:  A $d$-dimensional tensor $\mathcal{A}\in\mathbb{R}^{I_{1}\times\cdots\times I_{d}}$, fixed TT-ranks $\{r_{0},r_{1},\dots,r_{d}\}$ with $r_{0}=r_{d}=1$.
0:  Right-orthogonal TT-cores $\mathcal{G}^{(1)},\dots,\mathcal{G}^{(d)}$ of the approximation $\widehat{\mathcal{A}}$ with TT-ranks $r_{k}$’s.
1:  Temporary matrix: $\mathbf{C}=\mathrm{reshape}(\mathcal{A},\left[I_{1}\cdots I_{d-1},I_{d}\right])$, i.e., the $(d-1)$th unfolding matrix $\mathbf{A}_{d-1}$.
2:  for $k=d$ to $2$ do
@@
4:   Compute the truncated rank-$r_{k-1}$ approximation from URV decomposition $\mathbf{C}=\widehat{\mathbf{U}}\widehat{\mathbf{R}}\widehat{\mathbf{V}}^{\top}+\mathbf{E}$, where $\widehat{\mathbf{U}}\widehat{\mathbf{R}}\widehat{\mathbf{V}}^{\top}$ with $\widehat{\mathbf{R}}\in\mathbb{R}^{r_{k-1}\times r_{k-1}}$ is the leading term and $\mathbf{E}$ denotes the residual part in the URV decomposition (3).
5:   New core: $\mathcal{G}^{(k)}=\operatorname{reshape}\left(\widehat{\mathbf{V}}^{\top},\left[r_{k-1},I_{k},r_{k}\right]\right)$.
6:   $\mathbf{C}=\widehat{\mathbf{U}}\widehat{\mathbf{R}}$.
@@
8:  $\mathcal{G}^{(1)}=\mathbf{C}$.
Theorem III.2 (TT-URV case).
@@
Given a tensor $\mathcal{A}\in\mathbb{R}^{I_{1}\times\cdots\times I_{d}}$. Assume that the rank-$r_{k}$ approximation error of its $k$th unfolding matrix $\mathbf{A}_{k}\in\mathbb{R}^{(I_{1}\cdots I_{k})\times(I_{k+1}\cdots I_{d})}$ from the URV decomposition (3) is $\varepsilon_{k}$, that is,
$\mathbf{A}_{k}$ $=\begin{bmatrix}\mathbf{U}^{(k)}_{1}&\mathbf{U}^{(k)}_{2}\end{bmatrix}\begin{bmatrix}\mathbf{R}^{(k)}_{11}&\mathbf{R}^{(k)}_{12}\\ \mathbf{O}&\mathbf{R}^{(k)}_{22}\end{bmatrix}\begin{bmatrix}\mathbf{V}^{(k)\top}_{1}\\ \mathbf{V}^{(k)\top}_{2}\end{bmatrix}$
$=\mathbf{U}^{(k)}_{1}\mathbf{R}^{(k)}_{11}\mathbf{V}^{(k)\top}_{1}+\mathbf{E}^{(k)},$
@@
where $\mathbf{R}^{(k)}_{11}\in\mathbb{R}^{r_{k}\times r_{k}}$, $\mathbf{E}^{(k)}=(\mathbf{U}^{(k)}_{1}\mathbf{R}^{(k)}_{12}+\mathbf{U}^{(k)}_{2}\mathbf{R}^{(k)}_{22})\mathbf{V}^{(k)\top}_{2}$ and $\|\mathbf{E}^{(k)}\|=\varepsilon_{k}$ for all $k=1,\dots,d-1$. Then the approximate tensor $\widehat{\mathcal{A}}$ with TT-ranks not higher than $r_{k}$’s, computed via TT-URV algorithm 3 in the right-to-left sweep, meets (9)
Additionally, the TT-URV algorithm for prescribed accuracy employed in the right-to-left sweep can be obtained analogously. We will not repeat it here.
Remark III.2.
@@
Based on the above theoretical analysis, we recommend the TT-ULV algorithm in a left-to-right sweep to compute the left-orthogonal TT-cores, and the TT-URV algorithm in a right-to-left sweep to compute the right-orthogonal TT-cores, especially for the computations of the approximate TT tensor under a prescribed accuracy.
IV Numerical experiments
In this section, we demonstrate the good performances of the TT-UTV algorithms by means of different applications. There are a variety of available algorithms for rank-revealing UTV decomposition, such as the low-rank UTV algorithm [10] packaged in the ‘UTV tools’ [12, 11], which are computationally attractive alternatives to the SVD for low-rank problems. We used them in the following numerical experiments.
IV-A Discrete multivariate functions
The first example is the Hilbert tensor with elements
$\mathcal{X}_{i_{1}\cdots i_{d}}=\frac{1}{i_{1}+\cdots+i_{d}},$
which possesses low numerical TT-ranks structure [28]. We perform two groups of experiments to test the TT-UTV algorithms, one is on the third-order Hilbert tensor $\mathcal{X}$ of dimensions $160\times 160\times 160$, and the TT-ranks are taken as $(1,r,r,1)$ with $r=2,4,\cdots,30$. Figure 1 (a) shows the Frobenius norm errors of the TT approximations obtained by the TT-SVD algorithm in the left-to-right sweep, TT-UTV algorithms 1 and 3, respectively. Table I reports the concrete comparisons of the absolute errors of these algorithms under different TT-ranks, where the term ‘RtoL’ after TT-SVD means the algorithm is executed in the right-to-left sweep, and ‘LtoR’ stands for the left-to-right sweep. It can be seen that the TT decompositions computed by either TT-URV or TT-ULV provide satisfactory accuracy comparable to the TT-SVD.
Another is on the fourth-order Hilbert tensor of dimensions $50\times 50\times 50\times 50$, and its TT-ranks are taken as $(1,r,r,r,1)$ with $r=2,4,\cdots,20$, respectively. Figure 1 (b) shows that the errors decline with the increasing TT-ranks. The errors decrease exponentially as the rank increases, and the three curves are too close to distinguish. All these results show that the TT-UTV algorithms for computing the TT decomposition can provide the same accuracy as the TT-SVD algorithm.
Figure 1: The logarithms of absolute errors of the TT-approximations $\widehat{\mathcal{X}}$ obtained by various algorithms under different TT-ranks.
TABLE I: Absolute errors of the TT-approximations of the third-order Hilbert tensor
Fixed TT-ranks $(1,4,4,1)$ $(1,8,8,1)$ $(1,12,12,1)$ $(1,16,16,1)$ $(1,20,20,1)$
TT-SVD (LtoR) 3.43803418e-2 6.58860023e-5 7.37806779e-8 5.27161306e-11 1.40761816e-13
TT-ULV 3.43803409e-2 6.58860021e-5 7.37806778e-8 5.27161037e-11 1.24243097e-13
TT-SVD (RtoL) 3.43803418e-2 6.58860023e-5 7.37806779e-8 5.27161314e-11 1.40339130e-13
TT-URV 3.43803410e-2 6.58860021e-5 7.37806779e-8 5.27161611e-11 1.64036599e-13
Refer to captionRefer to captionRefer to caption
@@ IV-A Discrete multivariate functions
Refer to captionRefer to captionRefer to caption
(b) TT-SVD compressions
Refer to captionRefer to captionRefer to caption
(c) TT-ULV compressions
Refer to captionRefer to captionRefer to caption
(d) TT-URV compressions
Figure 2: Color images from the Berkeley Segmentation Dataset.
IV-B Color images
In this experiment, we test the performances of TT-UTV algorithms on image compression. Three color images are picked from the Berkeley Segmentation dataset [25]. In order to facilitate the dimension factorizations, these images are resized into $324\times 486$ dimensions [18] using bicubic interpolation through the MATLAB ‘imresize’ command. Each image is represented by three $324\times 486$ matrices that correspond to the three RGB channels, respectively. For each image, we firstly reshape the channel matrices into higher-order tensors of size $18\times 18\times 18\times 27$, then compute their low-rank approximation by TT-SVD and TT-UTV algorithms.
Figure 2 shows the performances of the TT-compressions on three images, where the TT-ranks are set as $(1,15,45,25,1)$, hence a TT-compressed image requires only $100035$ parameters, much lower than the original image of $472392$ parameters. As we can see, these compressed images are too similar to distinguish their resolution and retain most of the information in the original images. Figure 3 reports the relative square errors (RSEs) of the TT-format tensors for these images under different TT-ranks $(1,15,r,25,1)$ with $r=40,42,\cdots,60$. The RSEs of the three algorithms are very close.
Figure 3: The RSEs of the TT compressions of these color images decline with the TT-rank $r$.
IV-C MRI data recovery
When a portion of the information of high-dimensional data is missing or distorted, we can recover the data based on TT-format tensor completion, which can be solved by the Riemann gradient descent (RGrad) method [35]. Different from the traditional gradient descent method on Euclidean spaces, RGrad method uses the gradients on Riemannian manifolds formed by the fixed-rank TT-format tensors. From an initial TT-format tensor in the Riemannian manifold, the first step in each iteration is computing the Riemannian gradient $\mathcal{RG}(\mathcal{X}_{k})$ at the current point $\mathcal{X}_{k}$. Then compute the estimated tensor $\widehat{\mathcal{X}}_{k+1}=\mathcal{X}_{k}-\alpha_{k}\mathcal{RG}(\mathcal{X}_{k})$, where $\alpha_{k}$ is the step size, but it may not locate in the prescribed Riemannian manifold. Hence, the next step is retracting $\widehat{\mathcal{X}}_{k+1}$ back to the manifold by the TT-SVD algorithm to generate the next iteration point $\mathcal{X}_{k+1}$ [2, 42]. In this experiment, we explore the effectiveness of the RGrad method while the retracting step in each iteration is substituted by the TT-UTV algorithms.
(a) Original slice
(b) TT-SVD compression
(c) TT-ULV compression
(d) TT-URV compression
(e) The slice to be recovered
(f) RGrad method with TT-SVD
(g) RGrad method with TT-ULV
(h) RGrad method with TT-URV
Figure 4: The reconstruction of one slice $\mathcal{M}_{::36}$ of the MRI data.
The experimental results are evaluated by relative square error (RSE) and peak signal-to-noise ratio (PSNR), which are commonly used in the data inpainting tasks [24, 19]. Let ,^$\mathcal{M},\widehat{\mathcal{M}}\in\mathbb{R}^{I_{1}\times\cdots\times I_{d}}$ denote the ground-truth tensor and recovered tensor, respectively, then the PSNR is defined as
@@ IV-C MRI data recovery
We use the cubical MRI data  [^1] https://brainweb.bic.mni.mcgill.ca/brainweb/selection_normal.html, which is a third-order tensor denoted by $\mathcal{M}$ of size $181\times 217\times 181$. Firstly, we determine the truncated TT-ranks of this tensor for the Riemannian manifold. The TT-SVD and TT-UTV algorithms for the prescribed relative error $\varepsilon=0.1$ all produce the approximate tensors in the TT format with TT-ranks equal to $(1,55,41,1)$. We choose a slice $\mathcal{M}_{::36}$ from the MRI data shown in Figure 4, the four subfigures in the first line present the original slice and compressed ones via various TT decomposition algorithms. As we can see, these compressed slices preserve the most information in the original tensors.
@@
Afterwards, to test the effectiveness of the RGrad method with TT-UTV algorithms, the 70% elements of the original MRI data $\mathcal{M}\in\mathbb{R}^{181\times 217\times 181}$ are randomly removed and we would like to recover it. We perform three experiments by the RGrad method where the retracting steps are computed by TT-SVD, TT-ULV, and TT-URV algorithms, respectively. The second row in Figure 4 shows the original slice with missing pixels, and the recovered slices by RGrad methods with TT-SVD, TT-ULV, and TT-URV, respectively. These algorithms have good performances and recover most of the missing data of the original tensor. In addition, Figure 5 (a) reports that the RSEs of the recovered tensor decline with the iterations, which converge to a level close to $0.1$, and (b) shows the PSNRs increase in the process of restoration. These curves are too close to distinguish, which illustrates the RGrad method with TT-UTV is as effective as the RGrad method with TT-SVD.
@@
Figure 5: The RSEs decline and PSNRs increase with the iterations of the RGrad methods with various retracting algorithms.
V Conclusion
In summary, this letter contributes a novel approach called TT-UTV for tensor-train decomposition. It utilizes the virtues of UTV decomposition to compute the TT-format tensors; hence a variety of UTV algorithms can be employed, which require less computational cost than TT-SVD. Through rigorous error analysis, we clarify the effectiveness of these algorithms and the difference between the ULV and URV cases. We recommend the TT-ULV algorithms in a left-to-right sweep to compute the left-orthogonal cores, and the TT-URV algorithm in a right-to-left sweep to compute the right-orthogonal cores. We highlight the efficacy and versatility of the proposed algorithms on image processing and MRI data completion, which can provide comparable accuracy to the TT-SVD.
Appendix A Appendix

### v2

@@ 3.2 TT-URV case
0: A $d$th-order tensor $\mathcal{A}\in\mathbb{R}^{I_{1}\times\cdots\times I_{d}}$, and fixed TT-ranks $\{r_{0},r_{1},\dots,r_{d}\}$ with $r_{0}=r_{d}=1$.
0: Right-orthogonal $T$-cores $\mathcal{G}^{(1)},\dots,\mathcal{G}^{(d)}$ of the approximation $\widehat{\mathcal{A}}$ with TT-ranks $r_{k}$’s.
1: Temporary matrix: $\mathbf{C}=\mathrm{reshape}(\mathcal{A},\left[I_{1}\cdots I_{d-1},I_{d}\right])$, i.e., the (d1)(d-1)th unfolding matrix $\mathbf{A}_{d-1}$.
2: for $k=d$ to 22 do
@@
4:  Compute the truncated rank-$r_{k-1}$ approximation from URV decomposition $\mathbf{C}=\mathbf{U}_{1}\mathbf{R}_{11}\mathbf{V}_{11}^{\top}+\mathbf{E}$, where $\mathbf{U}_{1}\mathbf{R}_{11}\mathbf{V}_{1}^{\top}$ with $\mathbf{R}_{11}\in\mathbb{R}^{r_{k-1}\times r_{k-1}}$ is the leading term, and $\mathbf{E}$ denotes the residual part in the URV decomposition (6).
5:  New core: $\mathcal{G}^{(k)}=\operatorname{reshape}\left(\mathbf{V}_{1}^{\top},\left[r_{k-1},I_{k},r_{k}\right]\right)$.
6:  $\mathbf{C}=\mathbf{U}_{1}\mathbf{R}_{11}$.
@@
[…]
Theorem 3.2 (TT-URV case).
@@
Given a tensor $\mathcal{A}\in\mathbb{R}^{I_{1}\times\cdots\times I_{d}}$. Assume that the rank-$r_{k}$ approximation error of its $k$th unfolding matrix $\mathbf{A}_{k}\in\mathbb{R}^{(I_{1}\cdots I_{k})\times(I_{k+1}\cdots I_{d})}$ from the URV decomposition is $\varepsilon_{k}$, that is,
$\mathbf{A}_{k}=\begin{bmatrix}\mathbf{U}^{(k)}_{1}&\mathbf{U}^{(k)}_{2}\end{bmatrix}\begin{bmatrix}\mathbf{R}^{(k)}_{11}&\mathbf{R}^{(k)}_{12}\\ \mathbf{O}&\mathbf{R}^{(k)}_{22}\end{bmatrix}\begin{bmatrix}\mathbf{V}^{(k)\top}_{1}\\ \mathbf{V}^{(k)\top}_{2}\end{bmatrix}=\mathbf{U}^{(k)}_{1}\mathbf{R}^{(k)}_{11}\mathbf{V}^{(k)\top}_{1}+\mathbf{E}^{(k)},$ (12)
@@
where $\mathbf{R}^{(k)}_{11}\in\mathbb{R}^{r_{k}\times r_{k}}$, $\mathbf{E}^{(k)}=(\mathbf{U}^{(k)}_{1}\mathbf{R}^{(k)}_{12}+\mathbf{U}^{(k)}_{2}\mathbf{R}^{(k)}_{22})\mathbf{V}^{(k)\top}_{2}$, and $\|\mathbf{E}^{(k)}\|=\varepsilon_{k}$ are defined for all $k=1,\dots,d-1$. Then the approximate tensor $\widehat{\mathcal{A}}$ with TT-ranks not higher than $\{r_{k}\}_{0}^{d}$, computed via the TT-URV algorithm 3 in the right-to-left sweep, meets
$\|\mathcal{A}-\widehat{\mathcal{A}}\|_{F}\leq\sqrt{\sum_{k=1}^{d-1}\varepsilon_{k}^{2}}.$
The TT-URV algorithm for prescribed accuracy can be obtained analogously, referring to Algorithms 2 and 3. It is not repeated here.
@@
Based on the above theoretical analysis, we recommend the TT-ULV algorithm in a left-to-right sweep to compute the left-orthogonal TT-cores and the TT-URV algorithm in a right-to-left sweep to compute the right-orthogonal TT-cores.
Remark 3.2.
In practice, both the ULV and URV decompositions can be flexibly adapted for computing the TT decomposition with left-orthogonal or right-orthogonal TT-cores. Taking TT-URV as an example, while our theoretical analysis indicates that it is better suited for a right-to-left update pattern, one can still use it to compute left-orthogonal TT cores for a given tensor. This is achieved by simply reversing the index order of the input tensor and applying the TT-URV algorithm with a right-to-left sweep to update the computation on the reversed tensor.
3.3 Computational cost
In both the TT-SVD and TT-UTV algorithms, the predominant computational cost in each iteration arises from low-rank matrix approximation, implemented through either SVD or the more general UTV decomposition. This distinction also represents the essential difference between these two approaches. Consequently, comparing their computational complexities boils down to the differences in cost between matrix SVD and UTV decomposition.
The computational procedure for the SVD typically involves two key stages: an initial reduction of the matrix to bidiagonal form, which is a non-iterative process, followed by an iterative QR phase. For an $m\times n$ matrix where $m\geq n$, the SVD demands $O(mn^{2})$ floating-point operations [golub2013matrix]. In contrast, the algorithms for UTV decompositions encompass a family of related methods, including the high-rank UTV updating method [Stewart1992an], the low-rank revealing UTV decompositions [Fierro1997low], and the rank-revealing QR decomposition [CHAN198767, gu1996efficient], etc. Although rank-revealing UTV algorithms share the same asymptotic complexity of $O(mn^{2})$ with the SVD, the scaling coefficient for the dominant $mn^{2}$ term is generally lower [fierro1999utv, golub2013matrix]. This difference is intuitive, as a UTV decomposition imposes fewer structural constraints on its factors than an SVD, leading to potentially lower computational overhead. In addition, UTV algorithms often make better use of fast BLAS routines and are more flexible for parallel processing compared to the SVD, resulting in comparatively faster execution in practical applications [martinsson2020randomized, Martinsson2019rand].
The computational savings of UTV over SVD are amplified in the TT decomposition of a $d$th-order tensor $\mathcal{A}\in\mathbb{R}^{I_{1}\times I_{2}\times\cdots\times I_{d}}$. Since the TT decomposition requires d1d-1 sequential matrix factorizations, the savings achieved by UTV decomposition at each iteration accumulate across the entire hierarchy. As a result, the UTV-based methods enjoy a scaled computational advantage relative to the TT-SVD method, building upon their inherent efficiency at the single-matrix level.
4 Experiments
In this section, we demonstrate the good performance of the TT-UTV algorithms for different applications. There are various available algorithms for rank-revealing UTV decomposition, such as the low-rank UTV algorithm [Fierro1997low] packaged in the ‘UTV tools’ [fierro1999utv, fierro2005utv], which will be used in the numerical tests. The experiments were conducted on a computer equipped with an Apple M4 chip (featuring a 10-core CPU) and 24 GB of RAM, using MATLAB R2025a. In our benchmark studies, the TT-SVD algorithm is executed using the economy-size SVD command svd(A, ‘econ’) to ensure computational efficiency.
4.1 Video data
The experimental dataset was derived from a source video  [^1] https://youtu.be/X9cCgErwqQghttps://youtu.be/X9cCgErwqQg comprising 3,757 frames at a resolution of 2160×3840. Owing to the substantial memory requirements of the raw video, each frame was downsampled to 72×128 pixels using the MATLAB ‘imresize’ command. Hence, the subsequent separation of the RGB channels produced 11,271 images of size 72×128, which form a third-order tensor of dimensions 11271×72×128, denoted by $\mathcal{V}$. To benchmark the proposed algorithms, we evaluated their efficiency and accuracy on the tensor $\mathcal{V}$ under both fixed-rank and given-precision conditions. Let $\widehat{\mathcal{V}}$ denote the tensor reconstructed from the TT formats generated by these algorithms. The reconstruction accuracy is measured by the Relative Squared Error (RSE), defined as
$\mathrm{RSE}=\frac{\|\widehat{\mathcal{V}}-\mathcal{V}\|_{F}}{\|\mathcal{V}\|_{F}}.$
Computational efficiency is benchmarked by the runtime of each algorithm. To mitigate the effects of randomness, the reported runtime is the average elapsed time over 10 independent executions. Note that the TT-SVD and TT-ULV algorithms are executed via a left-to-right sweep. In contrast, the TT-URV algorithm is applied to the tensor with its indices reversed and proceeds via a right-to-left sweep. All of them yield left-orthogonal TT-core tensors.
The performance of these algorithms was evaluated under fixed TT-ranks $(1,R_{1},R_{2},1)$, as reported in Table 1. As we can see, the numerical results demonstrate that the proposed TT-UTV algorithms effectively reduce computational time while achieving satisfactory accuracy comparable to that of TT-SVD. Additionally, the UTV-based algorithms exhibit superior computational efficiency when smaller TT-ranks are required.
We also evaluated the fixed-precision variants of the algorithms under different error tolerance settings, with the error weights $w_{k}$ set uniformly to $1/\sqrt{2}$. As summarized in Table 2, the TT-ranks determined by the TT-UTV algorithms closely match those produced by the TT-SVD algorithm. As expected, larger error tolerances lead to smaller TT-ranks. In all cases, the actual computational errors remain strictly below, yet close to, the specified error tolerances $\varepsilon$, which confirms the effectiveness of our theoretical error bounds derived for the TT-UTV algorithm. Although the TT-ranks determined by the TT-UTV methods are occasionally slightly higher than those from TT-SVD, the TT-UTV variants often achieve higher accuracy as a compensatory benefit. Moreover, the UTV-based algorithms exhibit advantages in computational efficiency in low-rank scenarios.
Table 1: The numerical results under various fixed $T$-ranks.
    Time (s)     RSE
$(R_{1},R_{2})$     $(5,5)$ $(10,10)$ $(15,15)$ $(20,20)$     $(5,5)$ $(10,10)$ $(15,15)$ $(20,20)$
$T$-SVD     86.0 89.7 97.8 101.0     0.1179 0.0938 0.0832 0.0776
$T$-ULV     41.1 50.6 58.6 64.4     0.1179 0.0939 0.0835 0.0782
$T$-URV     39.2 51.5 58.3 64.8     0.1179 0.0939 0.0835 0.0781
Table 2: The numerical results of fixed precision algorihtms
      $\varepsilon=0.09$     $\varepsilon=0.1$     $\varepsilon=0.11$   Method     $(R_{1},R_{2})$ Time (s) RSE     $(R_{1},R_{2})$ Time (s) RSE     $(R_{1},R_{2})$ Time (s) RSE TT-SVD     $(38,6)$ 87.6 0.0885     $(27,4)$ 83.8 0.0995     $(19,3)$ 84.2 0.1064 TT-ULV     $(38,6)$ 86.4 0.0885     $(27,5)$ 69.4 0.0941     $(19,3)$ 59.8 0.1064 TT-URV     $(39,6)$ 87.7 0.0886     $(28,5)$ 71.0 0.0945     $(19,3)$ 60.3 0.1065
4.2 Image compression
In this experiment, we test the performance of TT-UTV algorithms on image compression. Three color images are picked from the Berkeley Segmentation dataset [martin2001database], which are of size $321\times 481$ pixels. In order to facilitate the dimension factorizations, they are resized into $324\times 486$ dimensions using bicubic interpolation via the MATLAB ‘imresize’ command [ko2020fast]. Then each processed image is represented by three $324\times 486$ matrices corresponding to its RGB channels. We first reshape the channel matrices into higher-order tensors of size $18\times 18\times 18\times 27$, and then compute their low-rank approximation using the TT-SVD and TT-UTV algorithms.
Fig. 2 demonstrates the compression performance on three test images under a fixed TT-rank of $(1,15,45,25,1)$. With this configuration, a compressed image requires only $100,035$ parameters, fewer than the $472,392$ parameters of the original. Visually, these compressed images are too similar to distinguish their resolutions and retain most of the structural and textural information from the original images. Furthermore, Fig. 3 reports the Relative Squared Errors (RSEs) under varying TT-ranks of the form $(1,15,r,25,1)$, where $r$ ranges from 40 to 50 by increments of 1. The results indicate that the RSEs achieved by the three algorithms are very close across all tested ranks.
@@ 4.2 Image compression
(b) $T$-SVD compressions
(c) $T$-URV compressions
(d) $T$-ULV compressions
Fig. 2: Three color images from the Berkeley Segmentation Dataset.
Fig. 3: The RSEs of $T$ compressions on these images are very close, and decline with the $T$-rank $r$.
4.3 MRI data completion
When a portion of the information in high-dimensional data is missing or distorted, we can recover the data based on TT-format tensor completion, which can be solved using the Riemannian gradient descent (RGrad) method [Steinlechner2016]. Unlike the traditional gradient descent method in Euclidean spaces, the RGrad method uses the gradients on Riemannian manifolds formed by the fixed-rank TT-format tensors. From an initial TT-format tensor in the Riemannian manifold, the first step in each iteration is computing the Riemannian gradient $\mathcal{RG}(\mathcal{X}_{k})$ at the current point $\mathcal{X}_{k}$. Then compute the estimated tensor $\widehat{\mathcal{X}}_{k+1}=\mathcal{X}_{k}-\alpha_{k}\mathcal{RG}(\mathcal{X}_{k})$, where $\alpha_{k}$ is the step size, but it may not locate in the prescribed Riemannian manifold. Hence, the next step is retracting $\widehat{\mathcal{X}}_{k+1}$ back to the manifold by TT-SVD algorithm to generate the next iteration point $\mathcal{X}_{k+1}$ [Cai2022provable, wang2019tensor]. In this experiment, we explore the effectiveness of the RGrad method while the retraction step in each iteration is substituted by the TT-UTV algorithms.
The experimental results are evaluated by Relative Square Error (RSE) and Peak Signal-to-Noise Ratio (PSNR), which are commonly used in data inpainting tasks [Li2017low, ko2020fast]. Let $\mathcal{M},\widehat{\mathcal{M}}\in\mathbb{R}^{I_{1}\times\cdots\times I_{d}}$ denote the ground-truth tensor and the recovered tensor, respectively. Then, the PSNR metric is defined as
@@ 4.3 MRI data completion
We use the cubical magnetic resonance imaging (MRI) data  [^2] https://brainweb.bic.mni.mcgill.ca/brainweb/selection_normal.htmlhttps://brainweb.bic.mni.mcgill.ca/brainweb/selection\_normal.html, which is a third-order tensor denoted by $\mathcal{M}$ of size $181\times 217\times 181$. Firstly, we determine the truncated TT-ranks of this tensor for the Riemannian manifold. We set the prescribed RSE $\varepsilon=0.1$ for the TT decomposition via the TT-SVD and TT-UTV algorithms; all of them produce the approximate tensors in the TT format with TT-ranks equal to $(1,55,41,1)$. We choose a slice $\mathcal{M}_{::36}$ from the MRI data shown in Fig. 4. The four subfigures in the first line present the original slice and the compressed ones via various TT decomposition algorithms. As we can see, these compressed slices preserve the most information from the original tensors.
(a) Original slice
(b) $T$-SVD compression
(c) $T$-ULV compression
(d) $T$-URV compression
(e) 70% missing pixels
(f) RGrad method with TT-SVD, PSNR = $30.0277$
(g) RGrad method with $T$-ULV, PSNR = $30.0279$
(h) RGrad method with $T$-URV, PSNR = $30.0274$
Fig. 4: The reconstruction of one slice $\mathcal{M}_{::36}$ of the MRI brain data.
@@
Afterwards, to test the effectiveness of the RGrad method with TT-UTV algorithms, 70% elements of the original MRI data $\mathcal{M}\in\mathbb{R}^{181\times 217\times 181}$ are randomly removed, and we would like to recover them. We perform three experiments using the RGrad method, where the retracting steps are computed by TT-SVD, TT-ULV and TT-URV algorithms, respectively. The second row in Fig. 4 shows the original slice with missing pixels and the recovered slices by RGrad methods with TT-SVD, TT-ULV and TT-URV, respectively. These algorithms have good performance and recover most of the missing data of the original tensor. In addition, Fig. 5 (a) reports that the RSEs of the recovered tensor decline with the iterations, which converge to a level of about the prescribed accuracy $0.1$, and (b) shows that the PSNRs increase in the process of restoration. These curves are too close to distinguish, which illustrates that the RGrad method with TT-UTV is as effective as the Rgrad method with TT-SVD.
(a) The RSEs of the recovered tensor $\widehat{\mathcal{M}}$
(b) The PSNRs of the recovered tensor $\widehat{\mathcal{M}}$
@@
Fig. 5: The RSEs decline and PSNRs increase with the iterations of the RGrad methods.
5 Conclusions
For the computation of tensor-train decomposition, we employed the rank-revealing UTV decomposition to compute low-rank matrix approximations within the computational process, leading to the TT-UTV algorithm. The algorithm provides a unified framework that incorporates a wide range of UTV-type decompositions for computing the TT decomposition, fully leveraging the advantages of UTV decompositions. Under the algorithmic framework, rigorous theoretical analysis establishes a tight error bound associated with the truncated error in each UTV approximation. Based on this analysis, distinct updating strategies are recommended for the ULV and URV variants, and the fixed-precision version of these algorithms are given. Numerical experiments across various application scenarios demonstrate the effectiveness of the proposed methods. For future work, integrating the UTV decomposition into parallel algorithmic frameworks for TT decomposition, along with the development of their randomized variants constitute an interesting research direction.
Appendix A